OpenStudy (anonymous):

easy explanation of this problem. . . . . of the 49 students in the club, 14 were girls. What was the ration of boys to girls in the club?

OpenStudy (anonymous):

14 girls 35 guys 35/14 = 5/2

OpenStudy (anonymous):

49-14=35 14 girls 35 boys 35/14 simplified is 5/2

OpenStudy (anonymous):

ratio of boys to girls is 5:2

OpenStudy (anonymous):

anther way girls = 14/49 = 2/7 and so boys( 7/7 - 5/7 = 2/7)

OpenStudy (anonymous):

I mean boys (7/7 - 2/7=5/7) so 5/7 of 49 is 35